# Diffusion-based Task-oriented Semantic Communications with Model Inversion Attack

### v1

@@ II-C System Performance Evaluation
In this paper, we assume that the receiver performs the image classification task, so that $\mathcal{R}$ can be trained using cross-entropy (CE) loss
$\mathcal{L}_{\text{CE}}(\phi,\theta)=-\sum_{i}p(y_{i})log(q(\hat{y}_{i})),$ (6)
where $p(y)$ and $q(y)$ represent the true probability distribution and predicted probability of $y$, and the classification accuracy $\gamma$ can be defined as the number of correctly classified images divided by the total number of input images.
Figure 2: Adversary prepares its dataset and accesses the transmitter repeatedly, with a well-trained strong classifier judging its outputs. $t$ is uniformly sampled from $[0,T]$. Green dash lines indicate the datasets that are used to train the strong classifier.
Generally, since $\mathcal{A}$ wants to recover the estimation of source images, PSNR and SSIM are commonly used in this scenario. PSNR is defined as
$\text{PSNR}=10\log_{10}\left(\frac{R^{2}}{\text{MSE}}\right),$ (7)
where $R$ is the maximum pixel value (e.g., 255 for 8-bit images), and MSE is the mean squared error. SSIM calculates the similarity between two images, expressed as
@@
where $\mu_{x}$ and $\mu_{\hat{x}}$ are the average pixel values of $x$ and $\hat{x}$, $\sigma_{x}^{2}$ and $\sigma_{\hat{x}}^{2}$ are the variances of $x$ and $\hat{x}$, $\sigma_{x{\hat{x}}}$ is the covariance between $x$ and $\hat{x}$, and $c_{1}$, $c_{2}$ are small constants that stabilize the division and avoid zero denominators. SSIM has a range of $[-1,1]$ and usually observed within the range $[0,1]$, where a larger value shows higher similarity between the two images.
In this paper, we give another index that measures the success rate of $\mathcal{A}$’s attack.
Definition (1): In task-oriented communication systems, the success rate of an attack is evaluated by the similarity between the task-relevant information in the attacker’s estimated image and that in the original image.
In our later experiments, we find that PSNR and SSIM may not be good indices in task-oriented system. PSNR and SSIM only indicate the similarity with original images in pixel-level. For a specific task, it is obvious that “task” information is more important than normal images, so focusing on how much task information (e.g., “label” in a category task) is recovered by $\mathcal{A}$ is more reliable, since we do not want $\mathcal{A}$ to acquire the purpose of the communication under the demand of privacy preserving. Details of implementation of Def. (1) are introduced in the next section.
@@ III Modules and Training
In this section, we will first introduce each module in our system model together with data flow, then we provide a detailed description of the training procedure based on our system design.
@@ III-A Encoder, Classifier and Adversary Modules
In related work [29], an auxiliary decoder $\mathcal{D}$ is commonly introduced to simulate an attacker during training, particularly in the context of adversarial learning frameworks. However, our approach does not adopt such attacker-aware training strategies. Instead, we define $\mathcal{A}$ as a model trained via model inversion attacks, solely for the purpose of evaluating the privacy risks after the transmission model is established. Importantly, $\mathcal{A}$ is not allowed to participate in the training or inference of the transmitter $\mathcal{T}$ and receiver $\mathcal{R}$. Likewise, if $\mathcal{D}$ is used in other methods, it should be restricted to the training phase and excluded from the evaluation phase. In our framework, since no adversarial training is involved, the auxiliary decoder $\mathcal{D}$ is entirely omitted.
To evaluate the semantic fidelity of adversary given Def. (1), we use a strong classifier $\bar{\Theta}$ that has more complex network structure than $\Theta$, to judge how much task-related information is contained in the image restored by the attacker $\mathcal{A}$, as shown in Fig. 2. It can be described by
$\hat{\dot{y}}=\bar{\Theta}_{\bar{\theta}}(\hat{x}),$ (9)
where $\bar{\theta}$ is the parameter set of $\bar{\Theta}$. Based on Def. (1), we make a hard decision on $\bar{\Theta}$’s output to get $\hat{\dot{y}}$ and compare it with $y$, which means we only select the category with the highest probability and ignore the soft information of other categories.
@@ III-B Diffusion Module
We use denoising diffusion implicit models (DDIM) [39] to train the diffusion module in our system. In the forward process, diffusion module gradually adds Gaussian noise to the extracted feature $f$ with a series of schedules $\alpha_{1},\dots,\alpha_{T}\in[0,1)$. To avoid ambiguity and facilitate discussion, in this section, let $f_{0}$ represent $f\in F$ mentioned earlier; that is, $f_{0}\sim p(f_{0})$ denotes the original feature, which is the output of encoder $\Phi$ and is not distorted yet. Then, noisy versions $f_{1},\dots,f_{T}$ are obtained by the following process in $T$ steps [32]:
@@
$p(f_{t}|f_{0})=\mathcal{N}\left(f_{t};\sqrt{\bar{\alpha}_{t}}f_{0},(1-\bar{\alpha}_{t}\mathbf{I}\right),\forall t\sim\mathcal{U}\left(\{1,\dots,T\}\right),$ (11)
@@
where $\mathcal{P}$ represents the power ratio between the signal component $\sqrt{\bar{\alpha}_{t}}f_{0}$ and noise component $\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t}$ obtained after partially adding noise, and varies with the values of $\alpha_{1}$, $\alpha_{T}$ and $T$. In our system, $f_{0}$ is power-normalized after generated by encoder $\Phi$. Hence, according to Eq. (14), it is equivalent to the signal $f_{0}$ passing through an additive Gaussian white noise (AWGN) channel with signal-to-noise ratio $\mathcal{P}$, which we refer to as “equivalent SNR” (ESNR). As an example, we marked the point $t=184$ in Fig. 3, where the energy of the signal component is approximately equal to that of the noise component, and the ESNR $p=0.01$ dB $\simeq 0$ dB.
@@
With a well-trained U-Net, the goal is to progressively denoise $f_{t}$ and to recover $f_{0}$ starting from a latent variable $f_{T}\sim\mathcal{N}(0,\mathbf{I})$. Specifically, on step $t-1$, given $f_{t}$ and $\epsilon_{t}$, we have
@@
To generate samples efficiently, DDIM allows to skip intermediate time steps. Instead of iterating over all $T$ steps, we can select a subset of time steps $\{t_{1},t_{2},\dots,t_{N}\}$ with $N\ll T$. This sub-sampling strategy reduces the number of denoising steps while maintaining high-quality generation, making it significantly faster than traditional diffusion models.
There is another problem that needs to be discussed when applying diffusion to communication system. Generally, diffusion module cannot exactly recover the corresponding estimation of the input feature, as they are designed to learn the overall distribution of $f_{0}$ rather than the precise structure of individual inputs. When sampling from random white noise, the generation process is guided by learned distribution $f_{0}\sim p(f_{0})$, and U-Net produces outputs that have the same distribution as $p(f_{0})$. But in communication systems, the transmitted information must be recovered in a specific order. To ensure the denoised signal meets this requirement, we can avoid adding excessive distortion to $f$ during the forward diffusion process, while proper guidance during the reverse process can further assist in generating outputs $\hat{f}$ that closely resemble the original $\hat{f}$.
Regarding the setting of $t$, let $0\leq T^{\prime}\leq T$, and perform
@@
in the self-noising module $\Psi$ before transmitting. In the receiver, $\Delta$ tries to recover $f_{0}$ using $T^{\prime}$ steps’ iteration that is the same as Eq. (17). This gives a way to control the weight of “task-relevant information” $\sqrt{\bar{\alpha}_{T^{\prime}}}f_{0}$ in the distorted output. The larger $T^{\prime}$ is, the less “task-relevant information” in $f_{T^{\prime}}$, and nearly no guidance involved when $T^{\prime}=T$ (in which diffusion module generates $f_{0}$ without specific order). Since Eq. (18) is valid for any $0\leq T^{\prime}\leq T$, we can train the diffusion module to denoise from $T$ and deploy the well-trained module that is set by $T^{\prime}$ when transmitting the signal, while receiver recovers the estimation of $f_{0}$ without other explicit guidance required. In this way, the choice of $T^{\prime}$ controls the distortion of the transmitter, hence affects the difficulty for adversaries to estimate the original images with task information in them.
To further enhance system performance, the receiver can utilize self-generated label guidance. Specifically, the classifier module $\Theta$ generates the output $\hat{y}$ bypassing the diffusion module $\Delta$ at first. The accuracy of $\hat{y}$ is higher when distortion on $f$ is lower or channel condition is favorable, which means that $\hat{y}$ could be more helpful in such cases. Then $\hat{y}$ can be used to guide the diffusion module $\Delta$ in the denoising process (Fig. 1). Different from calculating the accuracy of $\hat{y}$, here we use the soft version of $\hat{y}$ that all the category info in the vector will be utilized as guidance. Next, $\hat{z}$ and $\hat{y}$ are fed into $\Delta$ to produce $\hat{y}^{\prime}$, which is further reintroduced into $\Delta$ for refined denoising. By iterating this feedback loop multiple times, the accuracy of $\mathcal{R}$ is significantly improved. Consistent with guidance-based DDIM implementation, we provide explicit external control to $\Delta_{\delta}$ by modifying the noise estimation in Eq. (15) to
@@ III-C Diffusion Module Under Wireless Channel
Here we discuss the impact of wireless channel on the diffusion module in our system. To facilitate discussion, let $f_{T^{\prime}}=z^{\prime}$ and $h=1$, then receiver gets
@@
where $\epsilon_{\text{ch}}\sim\mathcal{N}(0,\sigma_{\text{ch}}^{2}\mathbf{I})$ represents channel noise. Obviously, the channel noise term adds extra distortion to $f_{T^{\prime}}$ and changes its distribution, thus affects the sampling process at the receiver. To tackle this issue, we introduce some noise margin for possible channel interference during the forward process.
@@
where $\epsilon_{T^{\prime\prime}}\sim\mathcal{N}(0,\mathbf{I})$. Note that the term $\sqrt{\bar{\alpha}_{T^{\prime}}}f_{0}$ remains the same as $f_{T^{\prime\prime}}$. For simplicity, let $\epsilon_{T^{\prime}}=\epsilon_{T^{\prime\prime}}=\epsilon$. Then, $\mathcal{T}$ transmits $f_{T^{\prime\prime}}$ to the wireless channel, and receiver obtains
@@
Now we need to find $T^{\prime\prime}$. Let $\gamma$ represent channel’s signal-noise power ratio, then we have
@@
where $D$ is the length of $\epsilon$ and $\mathbb{E}[\|\epsilon^{2}\|]=D$.
@@
If $h$ is involved (e.g., Rayleigh channel), then
$\hat{z}^{\prime\prime}=h\sqrt{\bar{\alpha}_{T^{\prime}}}f_{0}+h\sqrt{1-\bar{\alpha}_{T^{\prime\prime}}}\epsilon+\epsilon_{\text{ch}}.$ (26)
Let
$\epsilon_{\text{total}}=h\left(\sqrt{1-\bar{\alpha}_{T^{\prime\prime}}}\epsilon_{T^{\prime\prime}}-\sqrt{1-\bar{\alpha}_{T^{\prime}}}\epsilon_{T^{\prime}}\right)+\epsilon_{\text{ch}},$ (27)
and we can get
$\frac{\hat{z}^{\prime\prime}}{\hat{h}}=f_{T^{\prime}}+\frac{\epsilon_{\text{total}}}{\hat{h}}$ (28)
Receiver can perform the signal estimation using minimum mean square error (MMSE) once $\hat{h}$ is estimated. However, the discussion of $h$ here is beyond the scope of our paper. In this paper, we focus on improving the task execution performance when model inversion attack occurs due to neural-based nature of the semantic systems, and we use AWGN channel as an example to illustrate our points. We will leave the study of systems involving more complex channel conditions in the future work.
@@ III-D System Training Procedure
1:  Initialization: Initialize weights $W$ and biases $b$ in $\mathcal{T}$, $\mathcal{R}$, $\mathcal{A}$ and $\bar{\Theta}$.
@@
16:      $\bar{\Theta}_{\bar{\theta}}\left(cat(x,\dot{x},\hat{x},\ddot{x})\right)=\hat{\dot{y}}$.
17:      Use $\hat{x}$ to calculate MSE loss by Eq. (5); use $\hat{\dot{y}}$ to calculate CE loss by Eq. (6).
@@
19:  Output: Pre-trained network $\mathcal{T}$, $\mathcal{R}$, $\mathcal{A}$ and $\bar{\Theta}$.
Here we introduce the training procedure in our experiment. As shown in Fig. 1, $t$ is used as a global parameter that controls the distortion at the transmitter and helps the denoising at the receiver. All modules are neural-based except for self-noising module and wireless channel modules. As for wireless channel, we consider AWGN channel in the experiment. The self-noising module can be regarded as a channel compensator, ensuring that the signals received by the diffusion module appear as if they have passed through a Gaussian channel, thereby conforming to the input requirements for denoising. MNIST and CIFAR10 are used as datasets to train the system model. Given the small size of images and features in these datasets, we only employ a simple diffusion module that is easy to train and implement, as well as convenient for deployment considering the low-latency requirements of practical communication systems.
@@
1:  Initialization: Load pre-trained networks $\mathcal{T}$, $\mathcal{R}$ and $\mathcal{A}$.
@@
7:      $\Theta_{\theta}(cat(\hat{z},T^{\prime}))=\hat{y}$.
@@
13:           $\Theta_{\theta}(cat(\hat{f},\hat{z},T^{\prime}))=\hat{y}$.
@@
19:         $\Theta_{\theta}(cat(\hat{f},\hat{z},T^{\prime}))=\hat{y}$.
@@
For ease of tracking, we divide the training procedure into two parts. The first part consists of three stages, each designed to prepare a specific module, as outlined in Algorithm 1. We assume that when an adversary $\mathcal{A}$ attempts to “join” a transmitter-receiver pair, the transmitter $\mathcal{T}$ and receiver $\mathcal{R}$ have already been fully trained and are successfully deployed on the transmission nodes. Consequently, the training of $\mathcal{A}$ must occur after the completion of the $\mathcal{T}$$-\mathcal{R}$ pair. As shown in Fig. 2, we use the well-trained transmitter to train the adversary, where receiver is not involved here. Hence, the performance of $\mathcal{A}$ will inherently depend on the training outcomes of $\mathcal{T}$. By optimizing the training of the $\mathcal{T}$$-\mathcal{R}$ pair and subsequently training $\mathcal{A}$ to its fullest potential, we aim to establish a (non-strict) upper bound on the amount of valid information the adversary can extract under the given dataset and system. The training of strong classifier $\bar{\Theta}$ can be carried out simultaneously with $\mathcal{A}$, or after the training of $\mathcal{A}$ is finished.
The second part is fine-tuning the receiver given fixed $\Phi$ and $\Psi$, as described in Algorithm 2. We implemented DiffSem, which performs a single-pass pipeline processing on the receiver’s input, and an enhanced version named E-DiffSem, where “E” stands for “enhance”. E-DiffSem utilizes soft $\hat{y}$ as self-referential guidance to improve the performance of denoising. Specifically, self-referential $\hat{y}$ is iteratively utilized until the accuracy of $\hat{y}$ reaches stability. For both methods, the diffusion steps $N$ are set to $\min(T^{\prime},50)$. Experimental results and discussions will be elaborated in the next section.

### v2

@@ II-C System Performance Evaluation
Figure 2: The adversary prepares its dataset and accesses the transmitter repeatedly, with a well-trained strong classifier evaluating its outputs. Green dashed lines indicate the datasets that are used to train the strong classifier.
In this paper, we assume that the receiver performs an image classification task, so that $\mathcal{R}$ is trained using the cross-entropy (CE) loss
$\mathcal{L}_{\text{CE}}(\phi,\theta)=-\sum_{i}p(y_{i})\,\log q(\hat{y}_{i}),$ (6)
where $p(y)$ and q(y^)q($\hat${y}) denote the true and predicted label distributions, respectively. The classification accuracy is the fraction of correctly classified inputs.
Pixel-level evaluations. Since the adversary $\mathcal{A}_{a}$ aims to estimate source images, pixel metrics such as PSNR and SSIM are often employed:
$\text{PSNR}=10\log_{10}\!\left(\frac{R^{2}}{\text{MSE}}\right),\qquad$ (7)
@@
where $R$ is the maximum pixel value (e.g., 255 for 8-bit images), MSE is the mean squared error, $\mu_{x}$ and $\mu_{\hat{x}}$ are the average pixel values of $x$ and $\hat{x}$, $\sigma_{x}^{2}$ and $\sigma_{\hat{x}}^{2}$ are the variances of $x$ and $\hat{x}$, $\sigma_{x{\hat{x}}}$ is the covariance between $x$ and $\hat{x}$, and $c_{1}$, $c_{2}$ are small constants that stabilize the division and avoid zero denominators. Higher PSNR indicates better perceptual quality between the reconstructed and reference images. SSIM takes values in [1,1][-1,1] but typically lies within $[0,1]$ in practice, with larger values corresponding to higher structural similarity.
Strong classifier aided evaluation. Here we introduce a strong classifier $\bar{\Theta}_{\bar{\theta}}$, whose architecture is deliberately larger and more expressive than the receiver’s classifier $\Theta_{\theta}$, so that it can serve as a reliable module for evaluation. $\bar{\Theta}_{\bar{\theta}}$ is pretrained using training set including $x$, $\hat{x}$, and $\dot{x}$ and corresponding labels to categorize into the ground-truth label $\hat{\dot{y}}$ [^2] For simplicity, outputs of $\bar{\Theta}_{\bar{\theta}}$ are all denoted as $\hat{\dot{y}}$.. The strong classifier is never involved in training or inference of the system in Fig. 1; instead, it is only applied after transmission to assess the quality of the adversary’s reconstructions, as shown in Fig. 2. The role of $\bar{\Theta}_{\bar{\theta}}$ is to provide a reference that, if an attacker’s reconstruction still contains sufficient task-relevant semantics, the strong classifier is expected to classify them correctly. In this way, both the ordinary classifier $\Theta_{\theta}$ at the receiver side and the stronger evaluator $\bar{\Theta}_{\bar{\theta}}$ yield accuracies that can be directly compared, offering a clear measure of semantic fidelity.
@@ III Modules and Training
In this section, we first introduce each module in our system together with data flow, followed by a detailed description of the training procedure based on the system design.
@@ III-A Encoder, Classifier and Adversary Modules
In related work [29], an auxiliary decoder $\mathcal{D}$ is commonly introduced to simulate an attacker during training, particularly in the context of adversarial learning frameworks. However, our approach does not adopt such attacker-aware training strategies. Instead, we define $\mathcal{A}_{a}$ as a model trained via model inversion attacks, solely for the purpose of evaluating the privacy risks after the transmission model is established. Importantly, $\mathcal{A}_{a}$ is not allowed to participate in the training or inference of the transmitter $\mathcal{T}$ and receiver $\mathcal{R}$. Likewise, if $\mathcal{D}$ is used in other methods, it should be restricted to the training phase and excluded from the evaluation phase. In our framework, since no adversarial training is involved, the auxiliary decoder $\mathcal{D}$ is entirely omitted. The strong classifier $\bar{\Theta}_{\bar{\theta}}$, as mentioned above, is employed solely for evaluation of $\mathcal{A}_{a}$.
@@ III-B Diffusion Module
We use denoising diffusion implicit models (DDIM) [46] to train the diffusion module in our system. Unlike conventional generative usage, we adopt a paired design in which the transmitter deliberately applies a partial forward diffusion (self-noising) to $f_{0}$ up to a controllable step $T^{\prime}$, while the receiver runs a matching DDIM reverse schedule to deterministically denoise back toward $f_{0}$. This coupling makes $T^{\prime}$ an information knob that jointly governs utility and privacy. In the forward process, diffusion module gradually adds Gaussian noise to the extracted feature $f$ with a series of schedules $\alpha_{1},\dots,\alpha_{T}\in[0,1)$. To avoid ambiguity and facilitate discussion, in this section, let $f_{0}$ represent $f\in F$ mentioned earlier; that is, $f_{0}\sim p(f_{0})$ denotes the original feature, which is the output of encoder $\Phi_{\phi}$ and is not distorted yet. Then, noisy versions $f_{1},\dots,f_{T}$ are obtained by the following process in $T$ steps [33]:
@@
$p(f_{t}|f_{0})=\mathcal{N}\!\left(f_{t};\,\sqrt{\bar{\alpha}_{t}}\,f_{0},\,(1-\bar{\alpha}_{t})\,\mathbf{I}\right),\quad t\sim\mathcal{U}\{1,\dots,T\},$ (10)
@@
where $\mathcal{P}$ represents the power ratio between the signal component $\sqrt{\bar{\alpha}_{t}}f_{0}$ and noise component $\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t}$ obtained after partially adding noise, and varies with the values of $\alpha_{1}$, $\alpha_{T}$ and $T$. In our system, $f_{0}$ is power-normalized after generated by encoder $\Phi_{\phi}$. Hence, according to Eq. (13), it is equivalent to the signal $f_{0}$ passing through an AWGN channel with signal-to-noise ratio $\mathcal{P}$, which we refer to as “equivalent SNR” (ESNR). As an example, we marked the point $t=184$ in Fig. 3, where the energy of the signal component is approximately equal to that of the noise component, and the ESNR $p=0.01$ dB $\simeq 0$ dB.
Design rationale. The transmitter stops the forward diffusion at $T^{\prime}$ and the receiver runs the reverse schedule anchored at the same grid, so the receiver sees a corruption level consistent with its reverse schedule, which makes $T^{\prime}$ a trade-off parameter between system utility and attacker difficulty. In practice, $T^{\prime}$ is chosen once and reused across samples; DDIM’s snapped timestep grid further makes this setting robust to small SNR variations.
@@
With a well-trained U-Net, the goal is to progressively denoise $f_{t}$ and to recover $f_{0}$ starting from a latent variable $f_{T}\sim\mathcal{N}(0,\mathbf{I})$. Specifically, on step t1t-1, given $f_{t}$ and $\epsilon_{t}$, we have
@@
To generate samples efficiently, DDIM uses a deterministic, stride-based reverse schedule over a subset $\{t_{1}>\cdots>t_{N}\}\subseteq\{1,\dots,T\}$ with $N\!\ll\!T$. In our paired design, the receiver’s schedule snaps the target step $T^{\prime}$ to the nearest element in this grid.
A diffusion model learns the distribution $p(f_{0})$ rather than an exact per-sample inverse. When initialized from pure noise, it only ensures sampling from the learned distribution and thus cannot guarantee recovery of the specific input instance. However, communication demands recoverability of the transmitted semantics. Our paired design resolves this tension: the transmitter halts the forward diffusion early (self-noising at $T^{\prime}$) so that sufficient task-relevant structure remains in $f_{T^{\prime}}$, and the receiver runs the matching reverse schedule to deterministically peel off the injected noise. Optional guidance then further biases the reverse trajectory toward the intended task.
Transmitter-side self-noising. Let $0\leq T^{\prime}\leq T$, and perform
@@
in the self-noising module $\Psi$ before transmitting. In the receiver, $\Delta_{\delta}$ tries to recover $f_{0}$ using $T^{\prime}$ steps’ iteration that is the same as Eq. (16). This gives a way to control the weight of “task-relevant information” $\sqrt{\bar{\alpha}_{T^{\prime}}}f_{0}$ in the distorted output. The larger $T^{\prime}$ is, the less “task-relevant information” in $f_{T^{\prime}}$, and nearly no guidance involved when $T^{\prime}=T$ (in which diffusion module generates $f_{0}$ without specific order). Since Eq. (17) is valid for any $0\leq T^{\prime}\leq T$, we can train the diffusion module to denoise from $T$ and deploy the well-trained module that is set by $T^{\prime}$ when transmitting the signal, while receiver recovers the estimation of $f_{0}$ without other explicit guidance required. In this way, the choice of $T^{\prime}$ controls the distortion of the transmitter, hence affects the difficulty for adversaries to estimate the original images with task information in them.
To further enhance system performance, the receiver can utilize self-generated label guidance. Specifically, the classifier module $\Theta_{\theta}$ generates the output $\hat{y}$ bypassing the diffusion module $\Delta_{\delta}$ at first. The accuracy of $\hat{y}$ is higher when distortion on $f$ is lower or channel condition is favorable, which means that $\hat{y}$ could be more helpful in such cases. Then $\hat{y}$ can be used to guide the diffusion module $\Delta_{\delta}$ in the denoising process (Fig. 1). Different from calculating the accuracy of $\hat{y}$, here we use the soft version of $\hat{y}$ that all the category info in the vector will be utilized as guidance. Next, $\hat{z}$ and $\hat{y}$ are fed into $\Delta_{\delta}$ to produce $\hat{y}^{\prime}$, which is further reintroduced into $\Delta_{\delta}$ for refined denoising. By iterating this feedback loop multiple times, the accuracy of $\mathcal{R}$ is significantly improved. Consistent with guidance-based DDIM implementation, we provide explicit external control to $\Delta_{\delta}$ by modifying the noise estimation in Eq. (14) to
@@ III-C Diffusion Module Under Wireless Channel
Here we discuss the impact of wireless channel on the diffusion module in our system. To facilitate discussion, let $f_{T^{\prime}}=z^{\prime}$, then the receiver gets
@@
where $\epsilon_{\text{ch}}\sim\mathcal{N}(0,\sigma_{\text{ch}}^{2}\mathbf{I})$ represents channel noise. Obviously, the channel noise term adds extra distortion to $f_{T^{\prime}}$ and changes its distribution, thus affects the sampling process at the receiver. To tackle this issue, we introduce a noise margin for possible channel interference during the forward process.
@@
where $\epsilon_{T^{\prime\prime}}\sim\mathcal{N}(0,\mathbf{I})$. Note that the first term of $f_{T^{\prime\prime}}$ remains the same as that of $f_{T^{\prime}}$. For simplicity, let $\epsilon_{T^{\prime}}=\epsilon_{T^{\prime\prime}}=\epsilon$. Then, $\mathcal{T}$ transmits $f_{T^{\prime\prime}}$ to the wireless channel, and receiver obtains
@@
Now we need to find $T^{\prime\prime}$. Let $\gamma$ denote channel’s signal-to-noise power ratio (SNR), then we have
@@
where $D$ is the length of $\epsilon$ and $\mathbb{E}[\|\epsilon\|^{2}]=D$.
@@
Remark on SNR estimation (and sampler insensitivity). In the implementation, we simply substitute an estimate $\hat{\gamma}$ for $\gamma$ in Eq. (23)–(24). With a deterministic, stride-based DDIM sampler, the reverse process runs on a fixed discrete timestep sequence $\{t_{1}<\cdots<t_{N}\}\subseteq\{1,\dots,T\}$ (e.g., $t_{i}=10(i-1)$ for a 500-step DDPM schedule gives $0,10,20,\dots,500]$ under zero-based indexing). Consequently, the continuous target implied by Eq. (24) is effectively quantized to the nearest grid point when selecting $T^{\prime\prime}$. Small SNR mismatches typically unlikely to change the snapped index. In contrast, finely stepped stochastic DDPM schedules can be more sensitive to such perturbations. Hence, precise SNR estimation is not necessary for our method, as imprecise estimation only shifts the operating point marginally without affecting our conclusions.
@@ III-D System Training Procedure
1:Initialization: Initialize weights $W$ and biases $b$ in $\mathcal{T}$, $\mathcal{R}$, $\mathcal{A}_{a}$ and $\bar{\Theta}_{\bar{\theta}}$.
@@
16:    $\bar{\Theta}_{\bar{\theta}}\left([x;\dot{x};\hat{x};\ddot{x}]\right)=\hat{\dot{y}}$.
17:    Use $\hat{\dot{x}}$ to calculate MSE loss by Eq. (5); use $\hat{\dot{y}}$ to calculate CE loss by Eq. (6).
@@
19:Output: Pre-trained network $\mathcal{T}$, $\mathcal{R}$, $\mathcal{A}_{a}$ and $\bar{\Theta}_{\bar{\theta}}$.
Here we introduce the training procedure in our experiment. As shown in Fig. 1, $t$ is used as a global parameter that controls the distortion at the transmitter and helps the denoising at the receiver. All modules are neural-based except for self-noising module and wireless channel modules. The self-noising module can be regarded as a channel compensator, ensuring that the signals received by the diffusion module appear as if they have passed through a Gaussian channel, thereby conforming to the input requirements for denoising. MNIST and CIFAR-10 are used as datasets to train the system model. Given the small size of images and features in these datasets, we only employ a simple diffusion module that is easy to train and implement, as well as convenient for deployment considering the low-latency requirements of practical communication systems.
@@
1:Initialization: Load pre-trained networks $\mathcal{T}$, $\mathcal{R}$ and $\mathcal{A}_{a}$.
@@
7:    $\Theta_{\theta}(\hat{z},T^{\prime})=\hat{y}$.
@@
13:     $\Theta_{\theta}(\hat{f},\hat{z},T^{\prime})=\hat{y}$.
@@
19:     $\Theta_{\theta}(\hat{f},\hat{z},T^{\prime})=\hat{y}$.
@@
For ease of tracking, we divide the training procedure into two parts. The first part consists of three stages, each designed to prepare a specific module, as outlined in Algorithm 1. We assume that when an adversary $\mathcal{A}_{a}$ attempts to “join” a transmitter-receiver pair, the transmitter $\mathcal{T}$ and receiver $\mathcal{R}$ have already been fully trained and are successfully deployed on the transmission nodes. Consequently, the training of $\mathcal{A}_{a}$ must occur after the completion of the $\mathcal{T}$-$\mathcal{R}$ pair. As shown in Fig. 2, we use the well-trained transmitter to train the adversary, where receiver is not involved here. Hence, the performance of $\mathcal{A}_{a}$ will inherently depend on the training outcomes of $\mathcal{T}$. By optimizing the training of the $\mathcal{T}$-$\mathcal{R}$ pair and subsequently training $\mathcal{A}_{a}$ to its fullest potential, we aim to limit the amount of valid information that the adversary can extract under the given dataset and system. The training of strong classifier $\bar{\Theta}_{\bar{\theta}}$ can be carried out simultaneously with $\mathcal{A}_{a}$, or after the training of $\mathcal{A}_{a}$ is finished.
The second part is fine-tuning the receiver given fixed $\Phi_{\phi}$ and $\Psi$, as described in Algorithm 2. We implemented DiffSem, which performs a single-pass pipeline processing on the receiver’s input, and an enhanced version named E-DiffSem, where “E” stands for “enhance”. E-DiffSem utilizes soft $\hat{y}$ as self-referential guidance to improve the performance of denoising. Specifically, self-referential $\hat{y}$ is iteratively utilized until the accuracy of $\hat{y}$ reaches stability. For both methods, the diffusion steps $N$ are set to $\min(T^{\prime},50)$. Experimental results and discussions will be elaborated in the next section.